how do you find the slope of 4x+20y=5?
Found 2 solutions by rfer, edjones:
Answer by rfer(16322)

put in slope intercept for
y=mx+b
20y=-4x+5
y=-4/20x+5/20
y=-1/5x+1/4
slope =-1/5

Answer by edjones(8007)

y=mx+b Slope-intercept form. The slope is m.
4x+20y=5
20y=5-4x
y=5/20 - 4x/20
y=1/4 - x/5 The slope is -1/5.
.
Ed